You decide to make regular deposits every month into an account that earns 4% annual interest, compounded monthly. You hope to have $47000 at the end of 18 years. What is the amount of the regular payment you need to make?
Answer by Boreal(15235)

P=m(1+(r/n)^nk-1)/(r/n)
47000=m(1+.04/12))^216-1/(.04/12), multiply through by (0.04/12) and divide by the other term.
156.667=m(1.05) but round at the end
m=$148.93